The area of a rectangle is 60cm square. The length is 11cm more than the width. Find the width.
Answer by Boreal(15235)

width=w
length=w+11
w(w+11)=60
w^2+11w-60=0
(w-4)(w+15)=0
w=4 cm only positive root ANSWER
w+11=15, units are cm
the area is 15*4=60cm^2